OpenStudy (anonymous):

evaluate integral integral 1 to2 3/t^4 dt

OpenStudy (across):

\[\int_{1}^{2}\frac{3}{t^4}dt\]What's confusing you about this one?

OpenStudy (anonymous):

can you show me steps

OpenStudy (anonymous):

im not sure how i would solve this...would i have to reciprocal?

OpenStudy (anonymous):

\[\int_1^2{3t^{-4}}dt\]should be apparent now

OpenStudy (anonymous):

ok...how you go about doing that??

OpenStudy (anonymous):

first I take the constant 3 out\[3\int_1^2{t^{-3}dt}\]

OpenStudy (anonymous):

wouldnt that becomes 1/3?

OpenStudy (anonymous):

then I calculate the integral of \[t^{-3}dt\]before evaluating the definite integral:\[3\int_1^2{t^{-3}dt} = 3[\frac{t^{-2}}{-2}]_1^2 = -\frac{3}{2}[\frac{1}{t^2}]_1^2\]

OpenStudy (across):

\[\frac{a}{b}=ab^{-1}\]

OpenStudy (anonymous):

after that, you substitute the limits. \[-\frac{3}{2}[\frac{1}{4}-\frac{1}{1}]\]

OpenStudy (anonymous):

ok now i see i just dont understand the 3 constant taken out

OpenStudy (across):

Constants always get taken out of integrals (and differentials). For example, suppose I ask you to solve this:\[\int3dt.\]You can take the constant out:\[3\int dt,\]and still have the same expression:\[3t+C.\]

OpenStudy (across):

Also, it doesn't matter how tricky the constant looks:\[\int\frac{1}{3}dt=\frac{1}{3}\int dt,\]\[\int\pi dt=\pi\int dt.\]

OpenStudy (anonymous):

ok so in this case it would be just 3